OpenStudy (anonymous):

I don't understand the following relationship: Let M be a relation on integers such that (i, j) ∈ M if i ≡ j(mod 5) (0,0) ∈ M, (0,5) ∈ M, (5,0) ∈ M, (0,10) ∈ M (1,1) ∈ M, (1,6) ∈ M, (6,1) ∈ M, (1,11) ∈ M (2,2) ∈ M, (2,7) ∈ M, (7,2) ∈ M, (2,12) ∈ M (3,3) ∈ M, (3,8) ∈ M, (8,3) ∈ M, (3,13) ∈ M (4,4) ∈ M, (4,9) ∈ M, (9,4) ∈ M, (4,14) ∈ M It then says: [0]M = {...,−10,−5,0,5,10,15,...} [1]M = {...,−9,−4,1,6,11,16,...} [2]M = {...,−8,−3,2,7,12,17,...} [3]M = {...,−7,−2,3,8,13,18,...} [4]M = {...,−6,−1,4,9,14,19,...} It doesn't work though, for instance [1]_m =/= -9=-4(mod 5)

OpenStudy (amistre64):

looks like your getting into cosets, or partitions

OpenStudy (amistre64):

you need to get the mod number above 0, add another 5 to it

OpenStudy (amistre64):

-9 = 1 mod(5) -9 = [1]

OpenStudy (amistre64):

[1]M = {..., -9, -4, 1, 6, 11, 16, ...} ^ ^ -4 and 1 are both partitioned into the same set: -9 = [1] -4 = [1] 1 = [1]

OpenStudy (anonymous):

wait -9 = 1 mod(5)? I thought a relationship would be a Rb meant (a,b)∈ R therefore each numbers in the pair would be something like: -9= -4 (mod 5) should leave an answer of 1?

OpenStudy (anonymous):

but -9 = 1 ( mod 5) should give you a 2, no ? -9 - 1 / 5 ( unless I got this wrong)

OpenStudy (amistre64):

you have your set of integers partitioned into equivalence classes by convention, you (mod n) a value so that the remainder is between 0 and n -9 = a (mod 5) gives us a value of a that is between 0 and 5 -9 = 1 (mod 5) -4 = 1 (mod 5) 1 = 1 (mod 5) therefore -9 (mod 5) = -4 (mod 5) = 1 (mod 5)

OpenStudy (amistre64):

another way to look at it is to divide the integers like this ... -10 -9 -8 -7 -6 -5 -4 -3 -2 -1 0 1 2 3 4 5 6 7 8 9 ... each column is an equivalence class.

OpenStudy (anonymous):

I thought the modulo was the congruence of two numbers to their remainder ( right ? haha my teacher doesn't explain well) so -9 - 1 / 5 is congruent to 2 but -4 -1 /5 is congruent to 1. But if I look at the first set, the equivalence class is 0 as -5 = 0 ( mod 5) will give you a remainder of 0. Unless I'm seeing this all wrong.

OpenStudy (amistre64):

-9 = 5(-2) + 1

OpenStudy (amistre64):

you need to find an integer k, such that 5(k) is smaller than -9

OpenStudy (amistre64):

then the remainder defines the equivalence class by convention

OpenStudy (anonymous):

and in my equivalence class, will be defined by the remainder that any given pair will use right?

OpenStudy (amistre64):

the remainder will be in the same equivalence class. -9 = 5(-1) -4 ^^ remainder -9 = -4 they are in the same class, the same partition, but we have named the partitions as: [0] , [1] , [2] , [3] , [4] there is no [-4] partition so we need to establish what sort of remainders work for us, namely the remainders from 0 to 4

OpenStudy (anonymous):

Oh I see, the remainders aren't the equivalence class, but we decided to use them and we have to "fit" any other number included in our set in the equivalence class.

OpenStudy (amistre64):

since -9 = -4, where does -4 reside? -4 = 5(-1) + 1 ^^ -9 and -4 are in the partition [1]

OpenStudy (amistre64):

correct enough lol its best to find some integer multiple of 5, that is just smaller then the number we are modding for. negative number get smaller as the get 'bigger' :) -9 is bigger than -10 -9 = -10 + 1 -9 = 5(-2) + 1 -9 = 1 (mod 5)

OpenStudy (anonymous):

For that last example, you used k = -2 but that k resides in another equivalence class, are you allowed to do that?

OpenStudy (amistre64):

-4 = -5 + 1 hence: [1] = 5(k) + 1

OpenStudy (anonymous):

Im still trying to wrap my head around it but we're going for a = b ( mod 5 ) to a = 5(k) + 1 type thing, so I'm not seeing where the k comes from. Sorry for not understanding fast :/

OpenStudy (amistre64):

its an integer multiplier, and yes there is some connection to it that you ahvent gotten around to discussing yet. will any [2] equivalent work? -9 = 5(5(3)+2) + a ^^^^^ its in [2] -9 = 5(17) + a -9 - 5(17) = a -94 = a is -94 in [1] ? -94 = 5k + 1 -95 = 5k, it is :)

OpenStudy (amistre64):

you can show later on that mod partitions form groups, and groups can be algebraically manipulated.

OpenStudy (anonymous):

Oh I think I understand your point with the k. If I did -23 = 5k + 2, then 23 would be in the equivalence class of 2

OpenStudy (amistre64):

only if -23 - 2 is an integer multiple of 5 -25 = 5k is valid so yes

OpenStudy (amistre64):

-23 would be in the equivalance of [2]

OpenStudy (amistre64):

23 - 2 = 21, which isnt an integer multiple of 5 :) 23 is not in the equivalence of [2]

OpenStudy (anonymous):

So what does that have to do with the modulo. If I wrote -23 = 2(mod 5 ) , then the group (-23,2) would be in equivalence class 2?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

-23*2 = 2*2(mod 5) -26 = 4 (mod 5) (-26,4) is in [4]

OpenStudy (amistre64):

-46 not 26 .... i forgot how to multiply lol

OpenStudy (anonymous):

ahhh I see. But then, lastly, why are our equivalence class 0-4? Because those are not remainders, but more like integers to which if a number is added / subbtracted, it will be in the set of that number. So for instance Integer 5 will be in set 0 because 5 - 0 / 5 = 1, or am i seeing this wrong still

OpenStudy (amistre64):

we have to have some convention whereby we can organize the data. 5k + 0 5k + 1 5k + 2 5k + 3 5k + 4 these are all the possible remainders since k can be determined quite readily such that all integers fall into one of these 5 partitions

OpenStudy (amistre64):

are job is not to find some remainder that fits, but some k value that allows us to fit an integer into one of these 5 possible categories

OpenStudy (amistre64):

-9 = 5k + (0,1,2,3,4) -9 - (0,1,2,3,4) = 5k -9-1 allows us to determine k

OpenStudy (anonymous):

Oh I see... for that part haha. for me, remainders are ( like in programming) 5 % 4 = 1 because 4 * 1 + 1. so 5 - 4 = 1 or for instance 10 % 4 = 2, as 10 - 4(2) = 2 . In our case of b = a(mod 5) then a-b / 5 = 0,1,2,3 OR 4. If it equals to 0 ---> equivalence class of 0 for (a,b) and so on. This is what I'm understanding

OpenStudy (amistre64):

(a,b) in [n] if a = 5k + n and b = 5k + n for some n=0,1,2,3,4 -9 = 5(-2) + 1 -4 = 5(-1) + 1 (-9,-4) = [1]

OpenStudy (amistre64):

the value of k only needs to be some integer that makes it work

OpenStudy (amistre64):

im sure youve got the idea :)

OpenStudy (amistre64):

can you tell me if we partition the integers into groups of 8 93 = 8k + n, what is our equivalence class [n] going to be?

OpenStudy (anonymous):

Oh I see where youre getting at, it just confuses me as my notes use modulus and you us k. K being the number b and the number on the left being the number a, hence a = 5(b) + 1 a = 5(b) + 2 a = 5(b) + 3 a = 5(b) + 4 so is this right? and it will be... 5? as 93 % 8 = 5 , no?

OpenStudy (amistre64):

yes :) 93 is between 88 and 96 take the smaller value so we can 'add' n to it 93 = 88 + n n = 5, so 93 = 5 (mod 8) --------------------------------- (a,b) is not to be thought of as: a = 5b + 0 a = 5b + 1 a = 5b + 2 a = 5b + 3 a = 5b + 4 a and b are not related that way

OpenStudy (amistre64):

(a,b) in [n] if and only if a = 5k + n b = 5j + n for some integers k and j

OpenStudy (amistre64):

the gcd of a and b follows something similar to your idea, but we arent finding the gcd

OpenStudy (anonymous):

ah i see . thanks so much for your patience. id give you much more medals if i could !!!

OpenStudy (amistre64):

your welcome if youve covered gcd gcd (-9,-4) follows an algorithm -4 = 1(-9) + 5 1 = 0(5) + 1 <-- gcd = 1 0 = 1(0) + 0

OpenStudy (anonymous):

we did learn it, id have to go over it again, discrete math is complicated at first :/ thanks a lot :) ( again)

OpenStudy (amistre64):

not sure if that was wht you were confusing this with or not

OpenStudy (amistre64):

good luck :)